Shore Company reports the following information regarding its production cost. Units produced 42,000 units Direct labor $ 37 per unit Direct materials $ 38 per unit Variable overhead $ 294,000 in total Fixed overhead $ 108,920 in total Compute product cost per unit under absorption costing.
Answer #1

Variable overhead per unit=294,000/42,000=$7 per unit

Fixed overhead per unit=108,920/42,000=$2.59 per unit(Approx).

Product cost per unit under absorption costing=Direct materials+Direct labor+Variable overhead+Fixed overhead

=38+37+7+2.59

=$84.59 per unit(Approx).